# Learning Visual Generative Priors without Text

### v2

@@ 5 Conclusion
In this paper, we investigate image-to-image (I2I) generation, where models can learn from in-the-wild images in a self-supervised manner. First, we develop a purely vision-based training framework, Lumos, and validate the feasibility and scalability of learning I2I models. Our findings reveal that, as an upstream task of text-to-image (T2I) generation, our I2I model provides a more fundamental visual prior, achieving comparable or superior performance to current T2I models. It is worth mentioning that we only utilize 1/10 of the text-image pairs for fine-tuning of T2I models. Furthermore, we demonstrate the advantages of I2I priors over T2I priors in text-irrelevant visual generative tasks, such as image-to-3D and image-to-video generation. We believe this approach has the potential to liberate image generation from the constraints of image-text pairs, allowing it to learn a more foundational pure-visual prior.
@@ H.2 Lumos-T2I Model
1 256$\times$256 65 256x64 1.6x10-4 0.03 1
2 512$\times$512 60 64x64 8x10-5 0.03 1
3 1024$\times$1024 20 16x64 4x10-5 0.03 1
4 Multi-scale 1024 20 16x64 4x10-5 0.03 1
@@ I.4 Image Interpolation
  • “The picture shows a cute little tiger, wearing a blue hoodie and hat, sitting on a small cardboard boat on calm water..”

### v3

@@ 5 Conclusion
In this paper, we probe the image-to-image (I2I) generation, where models can learn from in-the-wild images in a self-supervised manner. First, we develop a purely vision-based training framework, Lumos, and validate the feasibility and scalability of learning I2I models. Our findings reveal that, as an upstream task of text-to-image (T2I) generation, our I2I model provides a more fundamental visual prior, achieving comparable or superior performance to current T2I models. It is worth mentioning that we only utilize 1/10 of the text-image pairs for fine-tuning of T2I models. Furthermore, we demonstrate the advantages of I2I priors over T2I priors in text-irrelevant visual generative tasks, such as image-to-3D and image-to-video generation. We believe this approach has the potential to liberate image generation from the constraints of image-text pairs, allowing it to learn a more foundational pure-visual prior.
Acknowledgment
This work is supported by National Nature Science Foundation of China (grant No.61871106), Key R&D projects of Liaoning Province, China (grant No.2024JH2/102500015). This work was supported by Ant Group Research Intern Program.
@@ H.2 Lumos-T2I Model
1 256$\times$256 65 256$\times$64 1.6$\times$10-4 0.03 1
2 512$\times$512 60 64$\times$64 8$\times$10-5 0.03 1
3 1024$\times$1024 20 16$\times$64 4$\times$10-5 0.03 1
4 Multi-scale 1024 20 16$\times$64 4$\times$10-5 0.03 1
@@ I.4 Image Interpolation
  • “The picture shows a cute little tiger, wearing a blue hoodie and hat, sitting on a small cardboard boat on calm water.”